OpenStudy (anonymous):

Use Euler’s Formula to find the missing number. Faces: 25 Vertices: 17 Edges: ?

OpenStudy (anonymous):

How do I use the Euler's Formula to find the answer?

OpenStudy (anonymous):

Do you know what Euler's formula for polyhedra is? \(V-E+F=2\)

OpenStudy (nathan917):

Easy. The formula is V-E+F=2. #1. F=25, V=17, so 42 - E = 2 E = 40

OpenStudy (anonymous):

Whew, that's what my brother said too. Thank you for explaining it @nathan917

OpenStudy (nathan917):

Np.